# Motion blur; Blind deblurring; Blind motion deblurring; Deep learning

### v1

@@ 2.4 Transformer-based Blind Motion Deblurring Methods
Synthetic Köhler at al. Köhler et al. (2012) Motion trajectory captured by 6D camera 4 sharp+48 blur Not divided
GoPro Nah et al. (2017) averaging over consecutive clear frames 3,214 pairs 2,103 / 0 / 1,111
HIDE Shen et al. (2019) averaging over consecutive clear frames 8,422 pairs 6,397 / 0 / 2,025
Blur-DVS Jiang et al. (2020) averaging over consecutive clear frames 2178 pairs 1782/ 0 / 396
Real RealBlur Rim et al. (2020)
@@
RsBlur Rim et al. (2022)
blurred images: The long exposure camera
clear images: the short exposure camera
@@
ReLoBlur Li et al. (2023)
blurred images: The long exposure camera
@@
2.5 Advantages and limitations of the four architectures.
In image deblurring, the four architectures CNN, RNN, GAN, and Transformer have their own advantages and limitations:
CNN is widely used in image processing to capture local features and spatial information Li et al. (2018); Liu et al. (2021). Trained with large-scale datasets, CNN-based deblurring methods perform high efficiency and generalization capabilities in Table 2. The CNN structure is simple and suitable for image denoising and deblurring tasks Tian et al. (2020). However CNN-based methods may be limited by a fixed-size receptive field and may not be effective for image-deblurring tasks that deal with global information or long-range dependencies. In order to solve the problem of a limited receptive field, the more widely used method is to use dilated convolution.
As shown in Table 2, RNN applies to sequence data and can capture temporal or sequential dependencies, which may be useful in certain image sequence deblurring scenarios. There may be gradient vanishing or explosion problems when dealing with long-term dependencies Rusch and Mishra (2021). Moreover, for image deblurring tasks, RNN is not well suited to capture spatial information Ren et al. (2021). Therefore, RNNs are generally combined with other structures to accomplish image deblurring tasks.
GAN improves the realism of image generation using adversarial training, which helps to generate more realistic deblurred images in Table 2. However, the training process may be unstable, requiring a balance between generator and discriminator training Cui et al. (2021). There may also be pattern crashes or non-convergence of training patterns.
As listed in Table 2, Transformer is able to capture global information and solve the remote spatial dependence problem Yan et al. (2023), and has advantages in processing for some image tasks that require long-distance dependence. However, the image deblurring task involves the processing of a large number of pixels and has a high computational cost.
3 Datasets and Evaluation Metrics
3.1 Datasets
Most neural network-based image deblurring methods typically require paired images for training. Existing blurry image datasets can be roughly divided into two categories: synthetic datasets and real datasets. Synthetic datasets include the Köhler dataset, Blur-DVS dataset, GoPro dataset, and HIDE dataset, among others. Real image datasets comprise the RealBlur dataset, RsBlur dataset, ReLoBlur dataset, and so on. As shown in Figure 9, these are example images from some commonly used datasets. Table 1 summarizes and outlines the basic information of select datasets. We additionally provide bar graphs to compare the size of the datasets, as shown in Figure 10.
Köhler et al. Dataset: To simulate real camera shake, Köhler et al. Köhler et al. (2012) used a Stewart platform with six degrees of freedom to replay the camera’s motion trajectory. During the replay of the motion trajectory, a series of clear images and long-exposure blurry images were captured. The dataset comprises 4 clear images and 48 corresponding blurry images for 12 motion trajectories.
@@ 3.1 Datasets
Blur-DVS Dataset: Created by Jiang et al. Jiang et al. (2020) using the DAVIS240C camera. The creation process involved capturing an image sequence with a slow-motion camera and then synthesizing 2,178 pairs of blurred and clear images by averaging adjacent 7 frames. The dataset consists of 1,782 pairs for training and 396 pairs for testing. Additionally, the dataset provides 740 real blurry images without paired clear images.
@@
GoPro Dataset: Developed by Nah et al. Nah et al. (2017), this dataset is one of the most commonly used deep learning methods. The dataset’s motion-blurred images are generated by integrating multiple instantaneous clear images within a time interval. Using a GoPro Hero4 Black camera capturing clear images at 240fps, they averaged over varying time windows to create blurred images, with the clear image at the window’s center serving as the ground truth. It includes 3,214 pairs of clear and blurred images, divided into 2,103 pairs for training and 1,111 pairs for testing.
@@
HIDE Dataset: Introduced by Shen et al. Shen et al. (2019), this dataset focuses on motion blur in pedestrian and street scenes, including camera shake and object movement. The dataset was created by capturing videos at 240 fps using a GoPro Hero camera and then averaging frames from these high-frame-rate videos. They averaged 11 consecutive frames to generate blurred images and kept the center frame as the clear image. It comprises 8,422 clear-blurred image pairs, divided into 6,397 pairs for training and 2,025 pairs for testing.
@@
RealBlur Dataset: Rim et al. Rim et al. (2020) created the RealBlur dataset to train and test deep deblurring methods on real blurry images. The dataset consists of two subsets sharing the same image content. One subset, RealBlur-J, is generated from the camera’s raw images, while the other, RealBlur-R, is generated from JPEG images processed by the camera’s ISP. It contains a total of 9,476 image pairs, with each subset offering 4,556 pairs of blurry and ground truth clear images for 232 low-light static scenes. The blurry images in the dataset are affected by camera shake and were captured in low-light scenarios such as night streets and indoor rooms, covering common motion blur scenes.
@@
RsBlur Dataset: Developed by Jaesung Rim et al. Rim et al. (2022), the RsBlur dataset provides real blurry images of various outdoor scenes, each paired with a series of 9 clear images. It comprises 13,358 real blurry images from 697 scenes. For analysis, the dataset is divided into training, validation, and testing sets, containing 8,878, 1,120, and 3,360 blurry images, respectively, with 465, 58, and 174 scenes in each set.
@@
ReLoBlur Dataset: Designed by Li et al. Li et al. (2023), ReLoBlur uses a simultaneous multi-band shutter device to capture local motion-blurred images along with their corresponding clear images, followed by post-processing for correction. Unlike synthetic blur datasets, ReLoBlur authentically showcases the mixing effect between locally moving objects and backgrounds, including overexposed, dimly lit real-life scenes, and complex blurred edges. It contains common real-life scenes such as pedestrians, vehicles, family members, pets, balls, plants, and furniture, among others. The dataset publicly provides 2,010 pairs of training set images and 395 pairs for testing. Additionally, the authors annotated the areas of local motion blur.
Figure 10: Histogram of the number of samples in the datasets.
Methods Pub. Category Overview Advantages Limitations
DeepDeblur Nah et al. (2017)
CVPR 2017
CNN
Removing complex motion blur from coarse to fine using a multi-scale recurrent network
End-to-end removal of complex motion blur without the need for kernel estimation
The mode has a large number of parameters,making training difficult
DMPHN Zhang et al. (2019)
CVPR 2019
CNN
Estimated by segmenting the image into multiple blocks and then removing blur at different scales
Overcoming the limitations of depth deblurring models on superposition depths
Feature segmentation may lead to discontinuous contextual information
PSS-NSC Gao et al. (2019)
CVPR 2019
CNN
Multi-scale CNN deblurring networks are analyzed and a parameter-selective sharing scheme is proposed
Selective parameter sharing can improve the learning efficiency and generalization of the model
The introduction of new network structures and hopping connection mechanisms may have increased computational costs
DGN Li et al. (2020)
TIP 2020
CNN
An effective depth-guided network consisting of an image deblurring branch and a depth refinement branch is proposed
The use of depth maps allows for a better understanding of scene changes and facilitates the use of depth information
Deblurring quality depends on the accuracy and quality of the depth map
MSCAN Wan et al. (2020)
TSCVT 2020
CNN
A multi-scale channel attention network is proposed for more robust feature representation and higher quality image recovery
Further taking advantage of multi-scale models, codec modules and residual block structures
The model has many parameters and is difficult to train
SDWNet Zou et al. (2021)
ICCV 2021
CNN
A network structure combining extended convolution and wavelet reconstruction is proposed for image deblurring
Dilated convolution results in a larger perceptual field and reduces the loss of texture information during sampling
Introduction of wavelet transform increases the complexity of the network
MIMOU-Net Cho et al. (2021)
ICCV 2021
CNN
Introducing asymmetric feature fusion for gradual recovery of clear images by coder-decoder
Handling multi-scale blurry with low computational complexity
Improvement of deblurring accuracy increases computational effort and time consuming
MPRNet Zamir et al. (2021)
CVPR 2021
CNN
Multi-stage architecture, injecting supervision at each stage to incrementally improve degraded inputs
Horizontal connections exist between feature processing blocks to avoid information loss
High dependence on computing power
MSSNet Kim et al. (2022)
ECCV 2022
CNN
A deep learning-based multi-scale stage network from coarse to fine is proposed
Multi-scale stage configurations and inter-scale information propagation contribute to the network’s deblurring performance
Hyperparameter adjustment is more complex
HINet Chen et al. (2021)
CVPRW 2022
CNN
Propose semi-instantiation to one piece using multi-stage architecture for image recovery task
HINet architecture generalizes well and has short inference time
Requires a large number of image pairs for training
BANet Tsai et al. (2022)
TIP 2022
CNN
Localization of blur direction and size by blurry perception for adaptive blur removal
Short inference time can well support subsequent real-time applications
Large amounts of diversity data are needed for training to adapt to dynamic scene blurring
IRNeXt Cui et al. (2023)
ICML 2023
CNN
Integration of multi-stage mechanisms into a u-shaped network to remove different sizes of blur in a coarse-to-fine manner
The local attention module improves restoration quality by weighting high-frequency information
The performance improvement of deblurring depends on the number of MSM branches
SVRNN Zhang et al. (2018)
CVPR 2018
RNN
End-to-end trainable spatially-variant RNN for dynamic scene deblurring
TThe pixel-by-pixel weights of the RNN are learned by the deep CNN, which helps to remove blurring
Need for a larger feeling field, dealing with large areas and spatial variations
SRN Tao et al. (2018)
CVPR 2018
RNN
Blur removal is achieved using a decoder with skip connections and parameter sharing across three scale
Sharing network weights at different scales, thus significantly reducing training complexity
Ignoring the variations between feature scales can easily saturate the network’s performance
MT-RNN Park et al. (2020)
ECCV 2019
RNN
Gradual removal of non-uniform blurring over multiple iterations using a self-looping MT structure
Contextual information is continuous, which facilitates the handling of blurring in dynamic scenes
Inflexible progressive training may not remove non-uniform blurring very well
DeblurGAN Kupyn et al. (2018)
CVPR 2018
GAN
Conditional GAN-based network for generating realistic deblurred images
Uniform and non-uniform motion blur images with less blurring can be recovered
Neglecting the nonlinearity characteristics of dynamic scenes blur, the network sensory field is small and homogeneous
DeblurGAN-V2 Kupyn et al. (2019)
ICCV 2019
GAN
Use feature pyramid networks and extensive backbone networks for better accuracy
Improved image restoration efficiency and enhanced real-time performance
Single network structure, which tends to lead to saturation of the network model
DBGAN Zhang et al. (2020)
CVPR 2020
GAN
Better deblurring by combining learning to generate blurred images and deblurring images
Synthetic blurred images more accurately mimic real blur for better deblurring
Network training is complex
CycleGAN Zhu et al. (2017)
ICCV 2021
GAN
Blind motion deblurring based on multi-adversarial mechanism for iterative generation of high-resolution images
Add multi-scale edge constraint function to enhance network deblurring and detail preservation capability
Presence of large color differences, severe artifacts and model redundancy
FCL-GAN Zhao et al. (2022)
MM 2022
GAN
Constraining Lightweight Network CycleGAN Driven Tasks by Frequency Domain Contrast Losses
Better deblurring of images in a lightweight and real-time manner
Deblurring effect does not reach the supervised model SOTA level
Ghost-DeblurGAN Liu et al. (2022)
IROS 2022
GAN
Introducing GhostNet for designing lightweight deblurring networks
Suitable for real-time image deblurring, small model size, fast inference time
Deblurring effect fails to reach SOTA level
Uformer Wang et al. (2022)
CVPR 2021
Transformer
Building hierarchical codec networks with the Transformer module
Ability to capture local and global dependencies
Self-attentive layers can only be applied to a limited depth
Restormer Zamir et al. (2022)
CVPR 2022
Transformer
Codecs for learning multi-scale local-global representations of high-resolution images
Ability to model global connectivity with high computational efficiency
Image recovery does not fully utilize spatial information
Stripformer Tsai et al. (2022)
ECCV 2022
Transformer
Re-weighting image features horizontally and vertically to capture blurred features in different directions
Low memory footprint and computational cost, does not rely on large amounts of training data
Scale dot products Note that complex matrix multiplication is usually required
FFTformer Kong et al. (2023)
CVPR 2023
Transformer
Combining a self-attentive solver with a feed-forward neural network based on discriminative frequency domain to remove blur
Lower space and time complexity and more effective and efficient in image deblurring
May not generalize well to different types of images and blurs
BiT Zhong et al. (2023)
CVPR 2023
Transformer
Introducing Double-Ended Temporal Supervision and Temporal Symmetry Integration Strategies to Generate Effective Time-Varying Motion Rendering Features
Capable of effectively handling blurring at different scales and fusing information from neighboring frames
Limited discrete supervision does not respond well to different realities
Sharpformer Yan et al. (2023)
TIP 2023
Transformer
Adaptive handling of non-uniform blurring with dynamic convolutional blocks by fully utilizing global-local features
Able to consistently improve the quality of deblurring due to CNN-based deblurring methods alone
Higher computational complexity
@@ 3.2.1 Subjective Evaluation
Subjective evaluation relies on human judgment and is often the most intuitive method for assessing image quality. One representative measure is the Mean Opinion Score (MOS) Hoßfeld et al. (2016). Subjective evaluation scales primarily include absolute and relative scales. As shown in Table 3, individuals rate image quality within the range of 1-5. However, when different deblurring methods produce similar results, subjective evaluation often struggles to differentiate subtle differences. This susceptibility to individual biases and external environmental factors during assessment can lead to varying subjective evaluation results among different observers. Therefore, subjective evaluation is generally used as a supplementary assessment method in image quality evaluation and cannot entirely determine the final image quality assessment outcome. As a result, for image deblurring, most existing methods rely on objective evaluation metrics for assessment.
@@
Score normal-$\uparrow$ Quality scale Relative quality Absolute quality
@@
Methods Category GoPro Nah et al. (2017) HIDE Shen et al. (2019) RealBlur-J Rim et al. (2020) Realblur-R Rim et al. (2020)
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
DeepDeblur Nah et al. (2017) CNN 29.08 0.914 25.73 0.874 27.87 0.834 32.51 0.841
DMPHN Zhang et al. (2019) CNN 31.20 0.940 29.09 0.924 28.42 0.860 35.70 0.948
SDWNet Zou et al. (2021) CNN 31.26 0.966 28.99 0.957 28.61 0.867 35.85 0.948
MSCAN Wan et al. (2020) CNN 31.24 0.945 29.63 0.920 - - - -
PSS-NSC Gao et al. (2019) CNN 31.58 0.948 - - - - - -
MIMOU-Net+ Cho et al. (2021) CNN 32.45 0.957 29.99 0.930 27.63 0.837 35.54 0.947
MPRNet Zamir et al. (2021) CNN 32.66 0.959 30.96 0.939 28.70 0.873 35.99 0.952
HINet Chen et al. (2021) CNN 32.77 0.959 - - - - - -
BANet Tsai et al. (2022) CNN 32.54 0.957 30.16 0.930 - - - -
SAPHNet Suin et al. (2020) CNN 32.02 0.953 29.98 0.930 - - - -
MSSNet Kim et al. (2022) CNN 33.01 0.961 30.79 0.939 28.79 0.879 35.93 0.953
LaKDNet Ruan et al. (2023) CNN 33.35 0.964 31.21 0.943 28.78 0.878 35.91 0.954
SRN Tao et al. (2018) RNN 30.26 0.934 28.36 0.915 28.56 0.867 35.66 0.947
MT-RNN Park et al. (2020) RNN 31.15 0.945 29.15 0.918 28.44 0.862 35.79 0.951
ESTRNN Zhong et al. (2023) RNN 31.07 0.902 - - - - - -
IFI-RNN Nah et al. (2019) RNN 29.97 0.895 - - - - - -
RNN-MBP Zhu et al. (2022) RNN 33.32 0.963 - - - - - -
DeblurGAN Kupyn et al. (2018) GAN 28.70 0.858 24.51 0.871 27.97 0.834 33.79 0.903
DeblurGAN-V2 Kupyn et al. (2019) GAN 29.55 0.934 26.61 0.875 28.70 0.866 35.26 0.944
DBGAN Zhang et al. (2020) GAN 31.10 0.942 28.94 0.915 24.93 0.745 33.78 0.909
CycleGAN Zhu et al. (2017) GAN 22.54 0.720 21.81 0.690 19.79 0.633 12.38 0.242
FCL-GAN Zhao et al. (2022) GAN 24.84 0.771 23.43 0.732 25.35 0.736 28.37 0.663
Ghost-DeblurGAN Liu et al. (2022) GAN 28.75 0.919 - - - - - -
CTMS Zhao et al. (2022) Transformer 32.73 0.959 31.05 0.940 27.18 0.883 - -
Uformer Wang et al. (2022) Transformer 32.97 0.967 30.83 0.952 29.06 0.884 36.22 0.957
Restormer Zamir et al. (2022) Transformer 32.92 0.961 31.22 0.942 28.96 0.879 36.19 0.957
Stripformer Tsai et al. (2022) Transformer 33.08 0.962 31.03 0.940 - - - -
Stoformer Xiao et al. (2022) Transformer 33.24 0.964 30.99 0.941 - - - -
FFTformer Kong et al. (2023) Transformer 34.21 0.969 31.62 0.945 - - - -
Sharpformer Yan et al. (2023) Transformer 33.10 0.963 - - - - - -
@@ 3.2.2 Objective Evaluation
Objective assessment primarily involves quantitatively analyzing and evaluating deblurred images through the construction of mathematical models Su et al. (2022); Li (2022); Wang et al. (2004). Objective assessment enables the unified measurement of various deblurring methods through precise data calculations, offering advantages of fairness, objectivity, and efficiency. Among these evaluation methods, Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) stand as the most commonly employed metrics in image restoration tasks.
Mean Square Error (MSE). MSE is the average of the squared absolute errors between the restored image and the ground-truth sharp image. The closer the MSE is to zero, the smaller the difference between the restored image and the ground truth image, indicating a more realistic restoration Li (2022).
@@
Peak Signal Noise Ratio (PSNR). PSNR Hore and Ziou (2010) is a commonly used metric to assess the effectiveness of image deblurring. It measures the accuracy of image deblurring by comparing the MSE between the deblurred and sharp images. However, PSNR only focuses on differences at the pixel level. In image deblurring assessment, it’s typical to combine other metrics to comprehensively evaluate image quality. PSNR is defined as:
@@
Structural Similarity Index Measure(SSIM). SSIM Wang et al. (2004) is one of the commonly used metrics in image quality assessment. In the task of image deblurring, it evaluates the similarity between the deblurred image and the ground-truth sharp image by comparing their structural information. SSIM primarily assesses image similarity based on three characteristics: luminance, contrast, and structure. It doesn’t just focus on pixel differences but also considers features like image structure and texture. Compared to PSNR, SSIM aligns better with human visual perception. The calculation formula is as follows:
@@
Table 5: Speed and Parameters of representative SOTA methods.
Method Category Param. (M) Time (s) PSNR (%)
DMPHN Zhang et al. (2019)
CNN
21.7
0.21
31.20
MIMOU-Net+ Cho et al. (2021)
CNN
16.1
0.02
32.45
MPRNet Zamir et al. (2021)
CNN
23.0
0.09
32.66
SRN Tao et al. (2018)
RNN
6.8
0.07
30.26
DeblurGAN-V2 Kupyn et al. (2019)
GAN
60.9
0.04
29.55
Restormer Zamir et al. (2022)
Transformer
26.1
0.08
32.92
Uformer Wang et al. (2022)
Transformer
50.9
0.07
32.97
Stripformer Tsai et al. (2022)
Transformer
19.7
0.04
33.08
FFTformer Kong et al. (2023)
Transformer
16.6
0.13
34.21

### v2

@@ 3.1 Datasets
Synthetic Köhler at al. [147] motion trajectory captured by 6D camera 4 sharp+48 blur Not divided
GoPro [79] averaging over consecutive clear frames 3,214 pairs 2,103 / 0 / 1,111
HIDE [148] averaging over consecutive clear frames 8,422 pairs 6,397 / 0 / 2,025
Blur-DVS [149] averaging over consecutive clear frames 2178 pairs 1782/ 0 / 396
Real RealBlur [150]
@@
RsBlur [151]
blurred images: long exposure camera
clear images: short exposure camera
@@
ReLoBlur [152]
blurred images: long exposure camera
@@
Blur-DVS Dataset: Created by Jiang et al. [149] using the DAVIS240C camera. The creation process involved capturing an image sequence with a slow-motion camera and then synthesizing 2,178 pairs of blurred and clear images by averaging adjacent 7 frames. The dataset consists of 1,782 pairs for training and 396 pairs for testing. Additionally, the dataset provides 740 real blurry images without paired clear images.
Figure 9: Histogram of the number of samples in the datasets.
@@
GoPro Dataset: Developed by Nah et al. [79], this dataset is one of the most commonly used deep learning methods. The dataset’s motion-blurred images are generated by integrating multiple instantaneous clear images within a time interval. Using a GoPro Hero4 Black camera capturing clear images at 240fps, they averaged over varying time windows to create blurred images, with the clear image at the window’s center serving as the ground truth. It includes 3,214 pairs of clear and blurred images, divided into 2,103 pairs for training and 1,111 pairs for testing.
@@
HIDE Dataset: Introduced by Shen et al. [148], this dataset focuses on motion blur in pedestrian and street scenes, including camera shake and object movement. The dataset was created by capturing videos at 240 fps using a GoPro Hero camera and then averaging frames from these high-frame-rate videos. They averaged 11 consecutive frames to generate blurred images and kept the center frame as the clear image. It comprises 8,422 clear-blurred image pairs, divided into 6,397 pairs for training and 2,025 pairs for testing.
@@
RealBlur Dataset: Rim et al. [150] created the RealBlur dataset to train and test deep deblurring methods on real blurry images. The dataset consists of two subsets sharing the same image content. One subset, RealBlur-J, is generated from the camera’s raw images, while the other, RealBlur-R, is generated from JPEG images processed by the camera’s ISP. It contains a total of 9,476 image pairs, with each subset offering 4,556 pairs of blurry and ground truth clear images for 232 low-light static scenes. The blurry images in the dataset are affected by camera shake and were captured in low-light scenarios such as night streets and indoor rooms, covering common motion blur scenes.
@@
RsBlur Dataset: Developed by Jaesung Rim et al. [151], the RsBlur dataset provides real blurry images of various outdoor scenes, each paired with a series of 9 clear images. It comprises 13,358 real blurry images from 697 scenes. For analysis, the dataset is divided into training, validation, and testing sets, containing 8,878, 1,120, and 3,360 blurry images, respectively, with 465, 58, and 174 scenes in each set.
@@
ReLoBlur Dataset: Designed by Li et al. [152], ReLoBlur uses a simultaneous multi-band shutter device to capture local motion-blurred images along with their corresponding clear images, followed by post-processing for correction. Unlike synthetic blur datasets, ReLoBlur authentically showcases the mixing effect between locally moving objects and backgrounds, including overexposed, dimly lit real-life scenes, and complex blurred edges. It contains common real-life scenes such as pedestrians, vehicles, family members, pets, balls, plants, and furniture, among others. The dataset publicly provides 2,010 pairs of training set images and 395 pairs for testing. Additionally, the authors annotated the areas of local motion blur.
@@ 3.2.1 Subjective Evaluation
Subjective evaluation relies on human judgment and is often the most intuitive method for assessing image quality. One representative measure is the Mean Opinion Score (MOS) [153]. Subjective evaluation scales primarily include absolute and relative scales. As shown in Table 3, individuals rate image quality within the range of 1-5. However, when different deblurring methods produce similar results, subjective evaluation often struggles to differentiate subtle differences. This susceptibility to individual biases and external environmental factors during assessment can lead to varying subjective evaluation results among different observers. Therefore, subjective evaluation is generally used as a supplementary assessment method in image quality evaluation and cannot entirely determine the final image quality assessment outcome. As a result, for image deblurring, most existing methods rely on objective evaluation metrics for assessment.
@@
Score $\uparrow$ Quality scale Relative quality Absolute quality
@@ 3.2.2 Objective Evaluation
Objective assessment primarily involves quantitatively analyzing and evaluating deblurred images through the construction of mathematical models [154, 155, 156]. Objective assessment enables the unified measurement of various deblurring methods through precise data calculations, offering advantages of fairness, objectivity, and efficiency. Among these evaluation methods, Mean Squared Error (MSE), Peak Signal-to-Noise Ratio (PSNR), and Structural Similarity Index (SSIM) stand as the most commonly employed metrics in image restoration tasks.
Mean Square Error (MSE). MSE is the average of the squared absolute errors between the restored image and the ground-truth sharp image. The closer the MSE is to zero, the smaller the difference between the restored image and the ground truth image, indicating a more realistic restoration [155].
@@
Peak Signal Noise Ratio (PSNR). PSNR [157] is a commonly used metric to assess the effectiveness of image deblurring. It measures the accuracy of image deblurring by comparing the MSE between the deblurred and sharp images. However, PSNR only focuses on differences at the pixel level. In image deblurring assessment, it’s typical to combine other metrics to comprehensively evaluate image quality. PSNR is defined as:
@@
Methods Category GoPro [79] HIDE [148] RealBlur-J [150] Realblur-R [150]
PSNR SSIM PSNR SSIM PSNR SSIM PSNR SSIM
DeepDeblur [79] CNN 29.08 0.914 25.73 0.874 27.87 0.834 32.51 0.841
DMPHN [87] CNN 31.20 0.940 29.09 0.924 28.42 0.860 35.70 0.948
SDWNet [142] CNN 31.26 0.966 28.99 0.957 28.61 0.867 35.85 0.948
MSCAN [141] CNN 31.24 0.945 29.63 0.920 - - - -
PSS-NSC [139] CNN 31.58 0.948 - - - - - -
MIMOU-Net+ [85] CNN 32.45 0.957 29.99 0.930 27.63 0.837 35.54 0.947
MPRNet [90] CNN 32.66 0.959 30.96 0.939 28.70 0.873 35.99 0.952
HINet [144] CNN 32.77 0.959 - - - - - -
BANet [145] CNN 32.54 0.957 30.16 0.930 - - - -
SAPHNet [88] CNN 32.02 0.953 29.98 0.930 - - - -
MSSNet [143] CNN 33.01 0.961 30.79 0.939 28.79 0.879 35.93 0.953
MSFS-Net [86] CNN 32.73 0.959 31.05 0.941 28.97 0.908 36.02 0.959
LaKDNet [158] CNN 33.35 0.964 31.21 0.943 28.78 0.878 35.91 0.954
MRLPFNet [84] CNN 33.50 0.965 31.63 0.946 - - - -
SRN [81] RNN 30.26 0.934 28.36 0.915 28.56 0.867 35.66 0.947
MT-RNN [91] RNN 31.15 0.945 29.15 0.918 28.44 0.862 35.79 0.951
ESTRNN [159] RNN 31.07 0.902 - - - - - -
IFI-RNN [160] RNN 29.97 0.895 - - - - - -
RNN-MBP [161] RNN 33.32 0.963 - - - - - -
DeblurGAN [99] GAN 28.70 0.858 24.51 0.871 27.97 0.834 33.79 0.903
DeblurGAN-V2 [103] GAN 29.55 0.934 26.61 0.875 28.70 0.866 35.26 0.944
DBGAN [106] GAN 31.10 0.942 28.94 0.915 24.93 0.745 33.78 0.909
CycleGAN [111] GAN 22.54 0.720 21.81 0.690 19.79 0.633 12.38 0.242
FCL-GAN [115] GAN 24.84 0.771 23.43 0.732 25.35 0.736 28.37 0.663
Ghost-GAN [108] GAN 28.75 0.919 - - - - - -
CTMS [54] Transformer 32.73 0.959 31.05 0.940 27.18 0.883 - -
Uformer [121] Transformer 32.97 0.967 30.83 0.952 29.06 0.884 36.22 0.957
Restormer [125] Transformer 32.92 0.961 31.22 0.942 28.96 0.879 36.19 0.957
Stripformer [124] Transformer 33.08 0.962 31.03 0.940 - - - -
Stoformer [162] Transformer 33.24 0.964 30.99 0.941 - - - -
FFTformer [126] Transformer 34.21 0.969 31.62 0.945 - - - -
Sharpformer [133] Transformer 33.10 0.963 - - - - - -
Structural Similarity Index Measure(SSIM). SSIM [156] is one of the commonly used metrics in image quality assessment. In the task of image deblurring, it evaluates the similarity between the deblurred image and the ground-truth sharp image by comparing their structural information. SSIM primarily assesses image similarity based on three characteristics: luminance, contrast, and structure. It doesn’t just focus on pixel differences but also considers features like image structure and texture. Compared to PSNR, SSIM aligns better with human visual perception. The calculation formula is as follows: